when a number x is subtracted from 36 and the difference is divided by x , the result is 2 . what is the value of x ?
(a) 12 (b) 17 (c) 18 (d) 21 (e) 24

Answer by mananth(16946)

36-x / x= 2
36-x=2x
3x=36
x=36/3
x=12